OpenStudy (anonymous):

Train A travels at 90 miles per hour and covers 360 miles. Train B covers the same distance but travels at 60 miles per hour. How much longer does it take Train B than Train A to cover that distance?

OpenStudy (anonymous):

How long does the train A take?

OpenStudy (anonymous):

90 miles per hour

OpenStudy (anonymous):

@sheika17 How long does the train A take?

OpenStudy (anonymous):

it take 90 miles per hour and covers 360 miles

OpenStudy (anonymous):

how many hours does train A covers 360 miles?

OpenStudy (anonymous):

S=V*T

OpenStudy (anonymous):

what does V mean and T mean?

OpenStudy (anonymous):

4 hours for the first train

OpenStudy (anonymous):

V means velocity

OpenStudy (anonymous):

T means time

OpenStudy (anonymous):

S means distance

OpenStudy (anonymous):

6 hours for the second one

OpenStudy (anonymous):

@eliassaab but how did you get that answer?

OpenStudy (anonymous):

\[ Time =\frac {360}{60} - \frac {360}{90}=6-4=2 \]

OpenStudy (anonymous):

'90 miles per hour' is the velocity velocityof A train ' 360 miles' is the distance that A covers

OpenStudy (anonymous):

@sheika17 did you get ti?

OpenStudy (anonymous):

\[ T=\frac D V=\frac{Distance}{Velocity} \]

OpenStudy (anonymous):

@eliassaab yes i did! you explained so much better. Thank you so much